# Investigation of the Impact of Synthetic Training Data in the Industrial Application of Terminal Strip Object Detection

### v1

@@ 3.5 Annotations
The comparison shows that including all visible parts of a terminal block into its bounding box inevitably leads to highly overlapping annotations, which is critical for most DL-based detection models. The major problem arises from the non-maximum suppression (NMS) algorithm commonly used in object detection. It is a post-processing step that iteratively selects the detection with the highest prediction score and removes all highly overlapping detections with the same predicted class label based on a threshold. In crowded scenes, NMS often results in a large number of false negatives, i.e., undetected objects, which compromises the overall detection performance [Jiang et al., 2023]. Therefore, we decided to restrict the overlaps of terminal block bounding boxes to 20% of the width of the thinner component, encouraging the DL models to learn tight positional predictions with reasonable overlaps during training. Additionally, we provide annotations for the end clamps, end covers, plug-in bridges, and test adapters to enable applications in which accessories should be detected.
@@ 4 EVALUATION
In this section, we evaluate the sim-to-real performance of standard object detection models in the industrial application of terminal strip object detection using the proposed image synthesis pipeline. Therefore, we selected the two standard model architectures, RetinaNet [Lin et al., 2017] and Faster R-CNN [Ren et al., 2017], representatively for the common single-stage and two-stage detectors and used different backbones according to the TensorFlow Object Detection (TFOD) framework [Huang et al., 2017]. As we found RetinaNet with a ResNet50 [He et al., 2016] backbone and Faster R-CNN with an Inception-ResNet [Szegedy et al., 2017] backbone to perform best in previous experiments without accessories, we leveraged these particular DL models for our present investigation of the domain gap.
The evaluation is based on the mean average precision (mAP) metrics mAP@0.5 and mAP@[0.5:0.95], as defined for the MS COCO [Lin et al., 2014] challenge. While the first considers a single intersection over union (IoU) threshold of 0.5, the second averages over ten equidistant IoU thresholds ranging from 0.5 to 0.95 to evaluate the predicted position of the objects of interest more strictly. Thus, it can be distinguished between an overall poor detection performance and correct detections with slightly inaccurate bounding boxes.
@@ 4.1 Datasets
For our experiments, we generated a synthetic dataset of 30,000 randomly configured terminal strip images at a resolution of $1024\times 512$ pixels using the image synthesis pipeline. It contains 36 frequently occurring terminal blocks in five different colors and suitable plug-in bridges, test adapters, end covers, and end clamps as additional accessories. The markings are only considered as distractors and not part of the detection, so 40 detection classes, composed of very similar and well-separable objects, are provided in total. The average number of objects per image is 25, whereas the accessories are overrepresented compared to the evenly distributed terminal blocks, as they occur in almost every image. The dataset is split into 79% train, 1% evaluation, and 20% test images to enable testing on synthetic images required to compare the performances of the DL models in simulation and reality.
To perform the real-world test, we used the first part of the image synthesis pipeline, as described in Subsection 3.1, to obtain 100 parts lists of random terminal strips and build them in reality. After that, we took three pictures of each terminal strip from different viewpoints and manually labeled the resulting 300 images for object detection. Hence, the image contents in the synthetic and the real-world datasets are very similar, enabling a meaningful investigation of the sim-to-real domain gap. Since some terminal blocks only marginally differ in their width, e.g., UT 4 and UT 6, we expect the DL models to confuse these classes in either domain.
@@ 4.2 Experiments
In our experiments, we leveraged the implementation of RetinaNet and Faster R-CNN available in the TensorFlow 2 Detection Zoo [^5] https://github.com/tensorflow/models/blob/master/research/object˙detection/g3doc/tf2˙detection˙zoo.md of TFOD. In this framework, the models can be configured via configuration files and fine-tuning from ImageNet [Deng et al., 2009] or MS COCO [Lin et al., 2014] pre-training is possible. Since we aim to provide the performance of standard DL models, we kept the default configuration except from the following changes:
@@
  • All data augmentation methods were removed since the dataset could be extended by generating more synthetic images if necessary, and further variations are also realizable in the image synthesis pipeline.
  • The aspect ratios of the anchor boxes were adapted to the most frequent bounding box ratios in the synthetic training dataset.
  • The number of training steps was set to 50,000.
  • The batch size for RetinaNet was set to eight and for Faster R-CNN to four due to hardware limitations.
  • ImageNet pre-training was selected.
Using these configurations, the RetinaNet and Faster R-CNN detection models were trained on the synthetic training data and evaluated on the synthetic and real test images to quantify the sim-to-real domain gap.
As we found the real images to show the terminal strips on a smaller scale than the synthetic training images, we further experimented with rescaling the real-world data. Therefore, we considered three different approaches to determine the rescaling factors. First, a constant factor of 1.5 was applied to all real images, as it seemed to scale the majority into the range of the synthetic images. Second, the image synthesis pipeline was used to generate 1,000 terminal strip images of various smaller scales annotated with the corresponding rescaling factors to fine-tune a simple ResNet50 regression model from ImageNet pre-training and predict the scaling factor for each real image separately. Finally, we considered the trained DL models as black-box functions and applied a Bayesian optimization algorithm to the real test images to obtain the best detection performance of RetinaNet and Faster R-CNN depending on the scaling factors. Accordingly, we refer to these experiments as real (1.5), real (ResNet), and real (Opt) in the following. Note that the latter approach is not applicable in a real-world application, as it requires ground-truth annotations. However, it demonstrates the potential performance when accurate predictions for image resizing are available and allows the evaluation of the ResNet50 regression model.
@@ 4.3 Results
The performances of RetinaNet and Faster R-CNN in our experiments are listed in Tables 1 and 2, respectively. As you can see, both models reach 99% $mAP@0.5$ and more than 92% $mAP@\left[0.50:0.95\right]$ on the synthetic test images, showing that terminal strip object detection is well-solvable in the simulated domain, and even very similar terminal blocks can be distinguished. Considering this as the reference performance, the sim-to-real domain gap seems to be huge in the case of unscaled real test images since the better Faster R-CNN only achieves 73.80% $mAP@0.5$, and the performance of both models drops by more than 20% when the predicted bounding boxes are evaluated more strictly.
Table 1: Comparison of the sim-to-real detection performances of RetinaNet based on different scaling factors.
Test dataset $mAP@0.5$ $mAP@\left[0.50:0.95\right]$
synthetic $99.26\%$ $92.57\%$
[…]
real (1.5) $70.02\%$ $46.42\%$
real (ResNet) $90.75\%$ $63.26\%$
real (Opt) $96.57\%$ $68.23\%$
Table 2: Comparison of the sim-to-real detection performances of Faster R-CNN based on different scaling factors.
Test dataset $mAP@0.5$ $mAP@\left[0.50:0.95\right]$
synthetic $99.08\%$ $94.81\%$
[…]
real (1.5) $88.79\%$ $61.11\%$
real (ResNet) $95.59\%$ $66.23\%$
real (Opt) $98.10\%$ $68.30\%$
@@
However, we find many terminal blocks to be misclassified as their smaller-sized variant, e.g., PT 4 as PT 2,5 or PT 4-TWIN as PT 2,5-TWIN, and that the number of undetected objects is high (see the appendix). Since this observation does not correspond to the results of the synthetic test and the real-world dataset tends to show the terminal strips on a smaller scale than the synthetic dataset, we assume that the size of the terminal strip components in the image is crucial to the distinction of similar objects that mainly differ in size. Hence, we briefly test this hypothesis by resizing all real images with a constant scaling factor of 1.5 and obtain a performance increase of more than 13% $mAP@0.5$ for both object detectors, as expected. Nevertheless, the sim-to-real performance gap is still significant in this experiment, and a more dynamic scaling approach would be required in practice anyway.
@@
One possibility for dynamically determining the scaling factor for a terminal strip image is to use a preprocessing DL model to predict it. Therefore, we again rendered synthetic images to train a ResNet50 regression model, accepting the additional influence of the domain gap on the detection performance. Utilizing this model to resize the real test images further improves the detection performances of both detectors significantly, reaching 90.75% and 95.59% $mAP@0.5$. This quantifies the domain gap to 8.51% for RetinaNet and 3.49% for Faster R-CNN, which is a remarkable result.
However, the potential of scaling factors has not yet been fully exploited, as demonstrated by the last experiment. Concretely, the mean absolute errors between the optimized and the predicted scaling factors are 0.11 for RetinaNet and 0.13 for Faster R-CNN, preventing the models from another 5.82% and 2.51% $mAP@0.5$ leap in performance. Thus, with optimized image resizing the sim-to-real domain gap reduces to 2.69% and 0.98% $mAP@0.5$ demonstrating the power of combining DR with domain knowledge even in challenging industrial applications.
@@
The main performance differences between synthetic and real test images seem to lie in the accuracy of the predicted bounding boxes, as the $mAP@\left[0.50:0.95\right]$ scores of RetinaNet and Faster R-CNN are more than 20% worse in comparison. Looking at the predictions, we observe that the white background often causes inaccurate bounding boxes when a terminal block is equipped with a marking at its upper or lower edge so that there is a low contrast between the object and background, as depicted in Figure 6. Furthermore, we belief that the manual annotations are not as exact as the calculated ones, introducing additional inaccuracies. Hence, the lower performance in $mAP@\left[0.50:0.95\right]$ is explainable and could be addressed in real-world applications, e.g., by ensuring backgrounds deviating in color.
@@ 5 DISCUSSION
The results show that a real-world industrial object detection task can be learned from synthetic training images based on already available CAD data. Although the Faster R-CNN seems more suitable for sim-to-real generalization, as it outperforms the RetinaNet in every experiment, the underlying Inception-ResNet consists of more than twice as many parameters as the ResNet50 backbone of RetinaNet, increasing the required hardware and computation time. However, both models perform well under optimized scaling conditions, qualifying this approach for potential use in industrial manufacturing. Additionally, we proved that suitable scaling factors can also be learned from simulated data and expect future work to further improve the image preprocessing.
Regardless, it depends on the requirements of the application if a detection rate of 98.10% $mAP@0.5$ is acceptable. For identifying already mounted terminal strip components in an electrical cabinet, e.g., this performance might be sufficient, whereas stricter regulations apply in quality control. In the latter case, however, simplified conditions such as a fixed camera position and knowledge about the background can be exploited, which we hypothesize has another positive impact on the performance, potentially satisfying these regulations. Hence, industrial requirements seem to be within the realm of possibility, although minor uncertainties of the expected performance remain.
Nevertheless, removing the intractability of collecting and annotating data facilitates the development of industrial vision systems that can solve more complex tasks than classical image processing methods and are much faster than manual inspections by humans. Even though it is not guaranteed that this approach satisfies the various industrial requirements in a specific use case, DL models could still assist manual processes saving costs and time. Therefore, the presented results should be considered as a promising baseline with room for improvements to encourage industrial manufacturers to contemplate image synthesis methods in their applications.
@@ 6 CONCLUSIONS
In this paper, we combined domain randomization and domain knowledge in an image synthesis pipeline to investigate the impact of synthetic training data in the industrial application of terminal strip object detection. Therefore, we trained the standard object detectors RetinaNet and Faster R-CNN on synthetic training images and evaluated their performance on a real-world test dataset. The results demonstrated the importance of scaling, specifically when some object classes only differ in size. Regardless, we showed that state-of-the-art image synthesis methods enable sim-to-real generalization with a manageable domain gap even in more complex industrial applications, removing the intractability of data collection that is the main obstacle for current DL-based vision tasks in the industry. Moreover, we provide a dataset of synthetic and real terminal strip images so other detection models and preprocessing methods can be benchmarked in this use case.
Since the performance of the DL models highly depends on image scaling in our experiments, in future work, we focus on learning a suitable preprocessing, including rotations, to reach the performance observed under optimized scaling conditions in a fully predictive manner. Moreover, additional object classes should be considered, as product portfolios of industrial manufacturers often consist of more than 40 components, to assess how the performance changes with the number of classes. Finally, we plan to investigate orthographic images of terminal strips since perspective images complicate the determination of accurate bounding boxes in the edge regions.
APPENDIX
@@
Randomly selected synthetic training images generated by the proposed image synthesis pipeline.
@@
Randomly selected images from the real-world dataset.
@@
Confusion matrix of the Faster R-CNN detection model evaluated on the unscaled real images at IoU and score thresholds of 0.5. While many correct detections can be seen on the diagonal, certain misclassifications often occur. Among them are terminal blocks such as PT 4 or UT 6 that are repeatedly misclassified as their smaller-sized variants. Although we expected these types of false predictions, they could not be observed in the synthetic tests, indicating a sim-to-real discrepancy in scaling.

### v2

@@ 3.5 Annotations
The comparison shows that including all visible parts of a terminal block into its bounding box inevitably leads to highly overlapping annotations, which is critical for most DL-based detection models. The major problem arises from the non-maximum suppression (NMS) algorithm commonly used in object detection. It is a post-processing step that iteratively selects the detection with the highest prediction score and removes all highly overlapping detections with the same predicted class label based on a threshold. In crowded scenes, NMS often results in a large number of false negatives, i.e., undetected objects, which compromises the overall detection performance [Jiang.2023]. Therefore, we decided to restrict the overlaps of terminal block bounding boxes to 20% of the width of the thinner component, encouraging the DL models to learn tight positional predictions with reasonable overlaps during training. Additionally, we provide annotations for the end clamps, end covers, plug-in bridges, and test adapters to enable applications in which accessories should be detected.
@@ 4 Evaluation
In this section, we evaluate the sim-to-real performance of standard object detection models in the industrial application of terminal strip object detection using the proposed image synthesis pipeline. Therefore, we selected the two standard model architectures, RetinaNet [Lin.2017] and Faster R-CNN [Ren.2015], representatively for the common single-stage and two-stage detectors, and the two more recent models YOLOv8 [yolov8_ultralytics] and DINO [zhang2022dino] to demonstrate robustness across different model architectures. For RetinaNet, Faster R-CNN, and DINO we used a ResNet50 [He.2016] backbone.
The evaluation is based on the mean average precision (mAP) metrics mAP@0.5 and mAP@[0.5:0.95], as defined for the MS COCO [Lin.2014] challenge. While the first considers a single intersection over union (IoU) threshold of 0.5, the second averages over ten equidistant IoU thresholds ranging from 0.5 to 0.95 to evaluate the predicted position of the objects of interest more strictly. Thus, it can be distinguished between an overall poor detection performance and correct detections with slightly inaccurate bounding boxes. We abbreviate mAP@[0.5:0.95] to mAP in the following for readability.
@@ 4.1 Datasets
For our experiments, we generated a synthetic dataset of 30,000 randomly configured terminal strip images at a resolution of $1024\times 512$ pixels using the proposed image synthesis pipeline. It contains 36 frequently occurring terminal blocks in five different colors and suitable plug-in bridges, test adapters, end covers, and end clamps as additional accessories. The markings are only considered as distractors and not part of the detection, so 40 detection classes, composed of very similar and well-separable objects, are provided in total. The average number of objects per image is 25, whereas the accessories are overrepresented compared to the evenly distributed terminal blocks, as they occur in almost every image. The dataset is split into 79% train, 1% evaluation, and 20% test images to enable testing on synthetic images required to compare the performances of the DL models in simulation and reality. Although a 1% evaluation split may appear small and unconventional, we found it to be sufficient as it still covers around 7,500 detection objects. Moreover, we did not intend to excessively tune hyperparameters.
To perform the real-world test, we used the first part of the image synthesis pipeline, as described in Subsection 3.1, to obtain 100 parts lists of random terminal strips and build them in reality. After that, we took three pictures of each terminal strip from different viewpoints and manually labeled the resulting 300 images for object detection. Hence, the image contents in the synthetic and the real-world datasets are very similar, enabling a meaningful investigation of the sim-to-real domain gap. This dataset is far more complex than the datasets in previous studies of the sim-to-real generalization performance and gives reasonable insights into the potential of synthetic training images in real-world industrial applications.
@@ 4.2 Experiments
In our experiments, we leveraged the model implementations from the Open MMLab Detection Toolbox [mmdetection] In this framework, the models can be configured via configuration files and several checkpoints, e.g., from MS COCO pre-training, are available. Since we aim to provide the performance of standard DL models, we kept the default configurations except from the following changes:
@@
  • All on-the-fly data augmentation methods were removed since the dataset could be extended by generating more synthetic images if necessary, and further variations are also realizable in the image synthesis pipeline.
  • If present, the aspect ratios of the anchor boxes were adapted to the most frequent bounding box ratios in the synthetic training dataset.
  • MS COCO pre-training was selected, even though we obtained similar results with ImageNet pre-training.
Using these configurations, the selected models were trained for 20 epochs with a batch size of 16 on a single NVIDIA A100 GPU with 80 GB of memory. To quantify the sim-to-real domain gap, we then evaluated the trained models on the synthetic test set and the manually labeled real images.
As we found the scale of the real images to have a high impact on the model performance, we further experimented with rescaling the real-world data. This sensitivity stems from the fact that standard CNNs are not equivariant to the dilatation group ($\mathbb{R}^{+}$, $\times$), meaning their performance degrades significantly under scale variations not seen during training. Therefore, we considered three different approaches to determine the rescaling factors for preprocessing. First, a constant factor of 1.5 was applied to all real images, as it seemed to scale the majority into the range of the synthetic images. Second, the image synthesis pipeline was used to generate 1,000 additional terminal strip images of various scales annotated with the corresponding rescaling factors to fine-tune a simple ResNet50 regression model from ImageNet pre-training and predict the scaling factor for each real image independently. Finally, we also explored the detection performance when scaling factors are irrelevant, such as in settings with a fixed camera position. To address this, we further employed Bayesian optimization with Gaussian Processes to automatically determine the optimal scaling factor for each real test image individually. In this formulation, the scaling factor was treated as a continuous optimization variable, and the objective function was defined as the F1-score of the detector on the given image. This approach serves two purposes: first, it provides an upper bound on the sim-to-real generalization performance when the effects of scale mismatch are mitigated, and second, it allows us to evaluate the ResNet50 regression model. We refer to these experiments as real (1.5), real (ResNet), and real (Opt) in the following.
@@ 4.3 Results
The performances of RetinaNet, Faster R-CNN, YOLOv8, and DINO in our experiments are listed in Table 1. All models reach 99% $mAP@0.5$ and more than 94% $mAP$ on the synthetic test images, showing that terminal strip object detection is well-solvable in the simulated domain, and even very similar terminal blocks can be distinguished. Considering this as the reference performance, the sim-to-real domain gap seems to be huge in the case of unscaled real test images since the best DINO model only achieves 87.30% $mAP@0.5$, and the performance drop is even more significant for the other models.
Table 1: Comparison of the sim-to-real detection performances of RetinaNet, Faster R-CNN, YOLOv8, and DINO based on different scaling methods for image preprocessing. The inference times were measured as the mean of 100 inference runs on random images of the real test dataset.
RetinaNet Faster R-CNN YOLOv8 DINO
Test dataset mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$
synthetic $4.30\% 99$$9.60\%$ $6.60\% 99$$9.40\%$ $7.30\% 99$$9.40\%$ $7.80\% 99$$9.90\%$
real $43.60\%$ $65.00\%$ $50.70\%$ $76.90\%$ $29.60\%$ $43.70\%$ $57.60\%$ $87.30\%$
real (1.5) $58.60\%$ $85.80\%$ $64.80\%$ $95.30\%$ $58.20\%$ $84.80\%$ $65.50\%$ $95.60\%$
real (ResNet) $63.60\%$ $93.00\%$ $64.60\%$ $96.00\%$ $63.00\%$ $92.40\%$ $65.90\%$ $97.40\%$
real (Opt) $66.70\%$ $97.70\%$ $66.30\%$ $98.30\%$ $66.80\%$ $97.80\%$ $66.70\%$ $98.40\%$
inference time $23.24ms$ $24.81ms$ $23.83ms$ $43.73ms$
@@
However, we find many terminal blocks to be misclassified as their smaller-sized variant, e.g., PT 4 as PT 2,5 or PT 4-TWIN as PT 2,5-TWIN (see Appendix A). Since this observation does not correspond to the results of the synthetic test and the real-world dataset tends to show the terminal strips on a smaller scale than the synthetic dataset, we assume that the size of the terminal strip components in the image is crucial to the distinction of similar objects that mainly differ in size. Hence, we briefly test this hypothesis by resizing all real images with a constant scaling factor of 1.5 and obtain the expected improvement with all models exceeding 84% $mAP@0.5$ and DINO even reaching 95.60%. While in other use cases the occurrence of different object scales could be addressed by varying object sizes during training data generation, this is no option when some object classes only differ in size. This necessitates an approach with only small scale variations in training and a dynamic scaling strategy during inference.
@@
One possibility for dynamically determining the scaling factor for a terminal strip image is to use a preprocessing DL model to predict it. Therefore, we again rendered synthetic images to train a ResNet50 regression model, accepting the additional influence of the domain gap on the detection performance. Utilizing this model to resize the real test images further improves the detection performances of all considered detectors with Faster R-CNN and DINO reaching 96.00% and 97.40% $mAP@0.5$ respectively. This quantifies the domain gap to 3.40% and 2.50% for these models at this point , which is a remarkable result.
However, the dependency of the performance on the object scale is very application-specific and does not transfer to any industrial parts detection application. To this end, we applied Bayesian optimization with Gaussian Processes to each model and each real test image individually to approach the detection performance when scaling is aligned in both domains. The results can be seen in the penultimate row of Table 1, where all models reveal a sim-to-real domain gap of less than 2% in $mAP@0.5$. This demonstrates the power of combining DR with domain knowledge even in the challenging industrial application of terminal strip object detection.
@@
The main performance differences between synthetic and real test images seem to lie in the accuracy of the predicted bounding boxes, as the mAP score does not exceed 66.80% in any experiment. Looking at the predictions, we observe that the white background often causes inaccurate bounding boxes when a terminal block is equipped with a marking at its upper or lower edge so that there is a low contrast between the object and background, as depicted in Figure 6. Furthermore, we belief that the manual annotations are not as exact as the calculated ones, introducing additional inaccuracies. Hence, the lower performance in mAP is explainable and could be addressed in real-world applications, e.g., by ensuring backgrounds deviating in color.
@@ 5 Discussion
Our studies show that real-world industrial object detection tasks come with their own characteristics and challenges that need to be addressed to introduce recent advances in DL to industrial environments. In particular, industrial settings often involve visually similar objects, repetitive arrangements, special object shapes, limited annotated data, and strict requirements for reliability and scalability. While previous research was more focused on the effect of individual image synthesis and randomization methods in simple domains rather than providing a usable pipeline for complex industrial applications, we developed a strategy that ensures a manageable implementation effort and demonstrated its strong baseline performance on this kind of applications.
However, there are a few things to consider to transfer our approach to a different industrial parts detection application. Assuming that detailed 3D models and expert knowledge about typical arrangements are available, it can be challenging to create realistic virtual materials that not only look like the real industrial parts in terms of color, but also have the same properties, e.g. reflection. This took us a few iterations as well, but one could also define a range of values and randomize the material during dataset generation to ensure robustness. Furthermore, the environment should be taken into account. In a more general case where the objects of interest can occur under various lighting conditions and in front of diverse backgrounds, it may be favorable to use a lot of different backgrounds and light settings in the training data, whereas under laboratory conditions the setups in reality and simulation can be aligned more closely to improve the performance.
In general, these advices together with the realistic arrangement of objects and the environment randomization already enable a cost effective generation of a synthetic image dataset for industrial parts detection. Nevertheless, application-specific issues can occur as we faced with scaling and object classes that only differ in size. The results show that these issues can harm the detection performance substantially. However, this gave us the chance to demonstrate that the same strategy can also be applied, e.g., to learn a suitable image preprocessing, even if our approach for predicting scaling factors leaves room for improvement. Regardless, we expect that this strategy can also help to overcome other application-specific issues without major changes.
As in limited data scenarios, data augmentation is another common way of extending the training dataset, the pros and cons of rendered data vs. augmented data should also be briefly discussed. While both approaches create synthetic images in a sense, a key difference is that data augmentation requires a base dataset of labeled images. On condition that no ground truth data of the target domain is available, which we focused on as a worst case consideration, this base dataset could only consist of synthetic images. Under this premise, data augmentation techniques such as variations in color, contrast or brightness could also be included in the simulation directly by changing the respective parameters in the rendering software. Hence, the only advantage of data augmentation would be that it is usually faster than rendering which could be relevant on a large scale. Regardless, augmentations must be selected as carefully as the randomization in the image synthesis pipeline to reflect domain-specific constraints. E.g. it would not make sense to randomly flip the images, as in our application terminal strips only occur in a single orientation. To support our assumptions on this, we performed an additional experiment in which we replaced $\frac{1}{3}$ of the original training data with some reasonable augmentations. The details and results can be found in Table 2 in the appendix.
It is important to note that we intend to provide a lower bound of the expectable detection performance in these complex industrial applications, which still can be improved. For instance, any ground truth data of the target domain is expected to be superior to synthetic data and to potentially further improve the performance. On that occasion, it would be interesting to benchmark augmented data of the target domain against synthetic data, which we leave for future work. To conclude, we think our results will serve as evidence for industrial manufacturers to start proof-of-concepts for their own use cases, as the cost-benefit ratio is promising.
@@ 6 Conclusions
In this paper, we investigated the impact of synthetic training data in the challenging industrial application of terminal strip object detection, which is representative for many detection tasks of industrial manufacturers. These tasks are often characterized by similar objects, repetitive arrangements, special object shapes, limited annotated data, and strict requirements for reliability and scalability discouraging companies from adopting DL solutions. By leveraging common randomization techniques and carefully incorporating domain knowledge into the image synthesis process, we demonstrated that it is possible to reach a detection performance suitable for industrial requirements with a manageable implementation effort. We also show that our approach is robust across model architectures, whereas the transformer based DINO model showed the best performance with 98.40% $mAP@0.5$ on our real-world test set.
We discussed the domain-specific issue of various scales in combination with objects that only differ in size, where we demonstrated that our strategy can be applied to learn a suitable image preprocessing from synthetic training data. This indicates that our approach is adaptable to various tasks in this context, although our learned image rescaling could not fix the issue entirely. We will further investigate this in future work. Moreover, additional object classes should be considered to scale up our approach, as product portfolios of industrial manufacturers often consist of more than 40 components, such as Phoenix Contact offering around 10,000 parts related to terminal strips. Finally, we plan to investigate orthographic images of terminal strips since perspective images complicate the determination of accurate bounding boxes in the edge regions.
In summary, our work establishes a practical and extensible strategy for leveraging fully synthetic data in complex industrial parts detection applications. It encourages manufacturers to consider image synthesis pipelines as a viable alternative to costly manual data collection and annotation, and provides a foundation for future advances in sim-to-real learning for industrial automation. In comparison to data augmentation, our approach has proven to be a strong alternative when no ground truth data of the target domain is available. Furthermore, we made our dataset publicly available for reference and future research to support further developments in this area.
Appendix A
@@
Figure 7: Randomly selected synthetic training images generated by the proposed image synthesis pipeline.
@@
Figure 8: Randomly selected images from the real-world dataset.
Table 2: Comparison of the sim-to-real detection performances of RetinaNet, Faster R-CNN, YOLOv8, and DINO based on different scaling methods for image preprocessing. For this experiment $\frac{1}{3}$ of the training data was replaced by data augmentation to investigate the impact of more generated data vs. augmented data. For dataset augmentation we used random brightness and contrast variations, Gaussian noise, motion blur, image compression, and slight perspective variations.
RetinaNet Faster R-CNN YOLOv8 DINO
Test dataset mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$ mAPmAP $mAP@0.5$
synthetic $93.30\%$ $98.80\%$ $6.10\% 99$$9.30\%$ $6.70\% 99$$9.30\%$ $7.40\% 99$$9.90\%$
real $39.10\%$ $58.70\%$ $47.80\%$ $71.90\%$ $22.80\%$ $34.80\%$ $55.60\%$ $86.20\%$
real (1.5) $59.00\%$ $86.30\%$ $64.80\%$ $94.70\%$ $56.90\%$ $86.70\%$ $65.50\%$ $98.10\%$
real (ResNet) $65.10\%$ $93.60\%$ $66.10\%$ $96.40\%$ $61.60\%$ $91.00\%$ $64.80\%$ $98.40\%$
real (Opt) $67.30\%$ $96.60\%$ $67.10\%$ $98.00\%$ $64.60\%$ $96.50\%$ $65.40\%$ $98.70\%$
@@
Figure 9: Confusion matrix of the Faster R-CNN detection model evaluated on the unscaled real images at IoU and score thresholds of 0.5. While many correct detections can be seen on the diagonal, certain misclassifications often occur. Among them are terminal blocks such as PT 4 or UT 6 that are repeatedly misclassified as their smaller-sized variants. Although we expected these types of false predictions, they could not be observed in the synthetic tests, indicating a sim-to-real discrepancy in scaling.